# Upper and Lower Bounds for the Linear Ordering Principle

### v1

@@ 6 Further Research
    One can define an input-oblivious version $\mathbf{{OL_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ of $\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ in different ways. One of the ways will provide a class that contains hard functions similarly to $\mathbf{{S_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ [Li23], $\mathbf{{O_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ [GLV24], and $\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ [KP24], and one can extend the proof of $\mathbf{{P}}^{\mathbf{{prMA}}}\subseteq\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ to show $\mathbf{{P}}^{\mathbf{{prOMA}}}\subseteq\mathbf{{P}}^{\mathbf{{OL_{2}^{\mathbf{{\scriptscriptstyle P}}}}}}$. We have not verified it so far and leave it as another conjecture for the moment. If true, it will also prove the Karp–Lipton–style theorem for $\mathbf{{OL_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$, which does not follow from known results otherwise.
@@ Acknowledgement
Appendix A A Karp–Lipton–style Collapse to $\mathbf{{P}}^{\mathbf{{prOMA}}}$
Proposition A.1.
Proof.
$\forall x\in\Pi_{\textsc{\tiny YES}}$ $\forall r\ \exists w\ A(x,r,w)=1,$
$\forall x\in\Pi_{\textsc{\tiny NO}}$ $\Pr_{r}[\exists w\ A(x,r,w)=1]<1/2.$
Let us ask Merlin to send a family $S_{n}$ of CircuitSAT circuits of appropriate input sizes (slightly abusing the notation: $S_{n}$ contains circuits for a polynomial range of input sizes, not just $n$); one can assume that they compute a correct satisfying assignment or say “no” (they can lie only if they say “no” for a satisfiable formula). Consider the Boolean circuit $A_{x,r}$ obtained by embedding $x$ and $r$ into $A$, its variables are the bits of the witness $w$. Now Arthur can use this circuit instead of Merlin’s original proof:
$\forall n\in\mathbb{N}\ \exists S_{n}\ \forall x\in\Pi_{\textsc{\tiny YES}}$ $\forall r\ A(x,r,S_{n}(A_{x,r}))=1,$
$\forall x\in\Pi_{\textsc{\tiny NO}}$ $\Pr_{r}[\exists S_{n}\ A(x,r,S_{n}(A_{x,r}))=1]<1/2.$
The first condition is true because Merlin can send correct CircuitSAT circuits. The second condition is true because if such a $S_{n}$ existed, then the original Merlin could have sent $w=S_{n}(A_{x,r})$. Note that in this case although $S_{n}$ may depend on $x$ and $r$, $A$ will able to catch a cheating Merlin. If we relax it by pulling the existential quantifier over $S_{n}$ outside, we get exactly the definition of $\mathbf{{prOMA}}$:
$\forall n\in\mathbb{N}\ \exists S_{n}\ \forall x\in\Pi_{\textsc{\tiny YES}}$ $\forall r\ A(x,r,S_{n}(A_{x,r}))=1,$
$\forall x\in\Pi_{\textsc{\tiny NO}}$ $\forall S_{n}\ \Pr_{r}[A(x,r,S_{n}(A_{x,r}))=1]<1/2.$
Appendix B Promise Oblivious Merlin–Arthur Protocols are in Promise $\mathbf{{O_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$
Proposition B.1.
Proof.
$\forall n\in\mathbb{N}\ \exists w_{n}\ \forall x\in\Pi_{\textsc{\tiny YES}}$ $\forall r\ A(x,r,w_{n}))=1,$ (1)
$\forall x\in\Pi_{\textsc{\tiny NO}}$ $\forall w\ \Pr_{r}[A(x,r,w)=1]<1/2.$ (2)
Condition (2) can be replaced by
Appendix C Loose Oracle Access vs Access Through a Consistent Language
Access through a consistent language.
The following definition is from [GS88].
Definition C.1 (promise problem, consistency).
Consider a promise problem $\Pi=(\Pi_{\textsc{\tiny YES}},\Pi_{\textsc{\tiny NO}})$, where $\Pi_{\textsc{\tiny YES}}\cap\Pi_{\textsc{\tiny NO}}=\emptyset$. We say that a language $O$ is consistent with $\Pi$ if $\Pi_{\textsc{\tiny YES}}\subseteq O$ and $\Pi_{\textsc{\tiny NO}}\subseteq\overline{O}$. The containment of $O$ outside of $\Pi_{\textsc{\tiny YES}}\cup\Pi_{\textsc{\tiny NO}}$ can be arbitrary. Let us denote it O
Π
fragments
[…]
Let us formalize it in the simplest case of $\mathbf{{P}}^{\Pi}$:
Definition C.2.
A language $L\in\mathbf{{P}}^{\Pi}$ according to the consistent language mode if there is a deterministic oracle Turing machine $M^{\bullet}$ with polynomial-time alarm clock stopping it in time $p(n)$ such that for every O
Π
fragments
[…]
and for every input $x$, $M^{O}(x)=L(x)$.
Definition C.3 (promise problems as oracles).
For more details and discussion see e.g. [GS88, BF99].
Loose access.
Let us define it rigorously in the simplest case of $\mathbf{{P}}^{\Pi}$:
Definition C.4.
A language $L\in\mathbf{{P}}^{\Pi}$ according to the loose access mode if there is a deterministic oracle Turing machine $M^{\bullet}$ with polynomial-time alarm clock stopping it in time $p(n)$ such that for every input $x$ $M^{\bullet}(x)$ produces the answer $L(x)$ if the oracle answers “yes” on queries in $\Pi_{Y}$ and answers “no” on queries in $\Pi_{N}$. (No hypothesis is made on its behaviour outside $\Pi_{Y}\cup\Pi_{N}$, in particular, its answers may be inconsistent for the same query.)
Equivalence for $\mathbf{{P}}^{\Pi}$.
Proposition C.5.
Proof.

### v2

@@ 6 Discussion and Further Research
    We defined an input-oblivious version $\mathbf{{OL_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ of $\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$, however, contrary to $\mathbf{{NP}}\subseteq\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$, it is not obvious whether $\mathbf{{ONP}}\subseteq\mathbf{{OL_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ or even $\mathbf{{ONP}}\subseteq\mathbf{{P}}^{\mathbf{{prOL_{2}^{\mathbf{{\scriptscriptstyle P}}}}}}$.
Relations between collapse results and “hard” functions.
Starting from Karp–Lipton’s paper [KL80], Kannan’s fixed-polynomial circuit complexity lower bounds [Kan82] were improving accordingly to new collapses: if a new collapse $\mathbf{{NP}}\subseteq\mathbf{{P/poly}}\implies\mathbf{{PH}}=\mathcal{C}$ is shown for a class $\mathcal{C}$ containing $\mathbf{{NP}}$, it immediately implies lower bounds for this class, because if $\mathbf{{NP}}\not\subseteq\mathbf{{P/poly}}$, we are already done.
However, a collapse to $\mathbf{{O_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ [CR06] did not imply lower bounds for $\mathbf{{O_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$, because $\mathbf{{NP}}$ is unlikely to be contained in it (after all, $\mathbf{{O_{2}^{\mathbf{{\scriptscriptstyle P}}}}}\subseteq\mathbf{{P/poly}}$). It was not until nearly two decades later that lower bounds for $\mathbf{{O_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ have been shown by Gajulapalli, Li, and Volkovich [GLV24] building on recent progress for the range avoidance problem [Kor22, Li24], thus matching the progress on the two questions again.
Korten and Pitassi [KP24] have shown fixed-polynomial lower bounds for their new class $\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ without showing a collapse result, thereby introducing a misalignment once again, yet this time in the opposite direction. Our paper’s inclusion $\mathbf{{P}}^{\mathbf{{prMA}}}\subseteq\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ restores the balance.
However, the observation that in the input-oblivious world the currently best collapse $\mathbf{{P}}^{\mathbf{{prOMA}}}$ reopens this question. Does this class possess fixed-polynomial circuit lower bounds? One can observe that Santhanam’s proof [San09] of $\mathsf{Size}[n^{k}]$ lower bounds for promise problems in $\mathbf{{prMA}}$ is input-oblivious. Indeed, the presented hard promise problems are actually in $\mathbf{{prOMA}}$! However, these promise problems do not yield a language in $\mathbf{{P}}^{\mathbf{{prOMA}}}$ that is hard for $\mathsf{Size}[n^{k}]$, and we leave this question for further research.
The best class for which fixed-polynomial circuit lower bounds can be proved (trivially) is $\mathbf{{P}}^{\textsc{$\varepsilon$\text{-}Hard-tt}}$ (for any particular fixed $\varepsilon>0$), where $\varepsilon$-Hard-tt, asks [^12] Korten [Kor22] defines a smoother version of it where the input length is not necessarily a power of two. given $1^{2^{n}}$, to output a truth table of a function $\{0,1\}^{n}\to\{0,1\}$ of circuit complexity at least $2^{\varepsilon n}$. Can one prove a collapse to this class (or at least to $\mathbf{{P}}^{\textsc{Avoid}}$)? Note that for the purpose of fixed-polynomial lower bounds even a limited version of $\varepsilon$-Hard-tt suffices where the truth table is non-empty for a number of entries greater than any polynomial and its complexity is only superpolynomial.